OpenStudy (anonymous):

A cylinder with radius r and height 2r + 4 contains a cube with edge length r√(2), as shown. What fraction of the cylinder's volume is taken up by the cube? Write your answer in simplified form.

OpenStudy (anonymous):

so 1st show the figure? :P

OpenStudy (anonymous):

Do you need it? @luckythebest

OpenStudy (anonymous):

or wait not really. volume of cylinder = Pi x r^2 x h => Pi x r^2 x (2r+4) cm3. Side of cube = r x root 2 volume = s^3 = r^3 x (root 2)^3 => r^3 x 2 x root 2 => \[2r^3\sqrt{2}\] cm3

OpenStudy (anonymous):

% of volume taken = \[\frac{ 2r^3\sqrt{2} }{ \Pi . r^2 . 2(r+2) }\]

OpenStudy (anonymous):

Here is the pic anyways:

OpenStudy (anonymous):

= \[=> \frac{ r \sqrt{2} }{ \Pi. (r+2) } x 100 \]

OpenStudy (anonymous):

I think this is a rational function or expression, which have equations of the form of f(x) = polynomial/polynomial. The graph of a rational function may have vertical and horizontal asymptotes.

OpenStudy (anonymous):

I didn't understand what you did that much.

OpenStudy (nubeer):

first find the volume of cylinder.. can you do that?

OpenStudy (anonymous):

Okay I will move on to another. Can you help me with this one? In a bike race, a rider covers a 5-mi flat stretch of road at a speed of s mi/h. She then doubles her speed down a hill 1 mi long. Finally, she reduces her downhill speed by 12 mi/h as she rides the last 3 mi of the race. what functions gives the time t it takes the rider to finish the reace in terms of ? and no.

OpenStudy (anonymous):

Did @luckythebest do it wrong @nubeer?

OpenStudy (nubeer):

no he did fine..

OpenStudy (anonymous):

Can you help me? @nubeer

OpenStudy (anonymous):

Hello? @nubeer

OpenStudy (nubeer):

umm sorry i thought @luckythebest was helping u here..

OpenStudy (nubeer):

time = distance * speed.. now we look at data. first 5 miles with s speed t1 = 5s then double speed for 1 mile t2 = 2s then speed reduced by 12 afor 3 miles t3= (2s-12) * 3 now add all t's

OpenStudy (nubeer):

@Mama123 ..

OpenStudy (anonymous):

how? @nubeer

OpenStudy (anonymous):

can you help or not? @luckythebest

OpenStudy (nubeer):

don't u know how to do addition.

OpenStudy (anonymous):

yes. is this right?: t^3 + 42 = S. @nubeer

OpenStudy (nubeer):

how did u got this?

OpenStudy (anonymous):

tell me if it is right or wrong first. then i will tell you how i got it.

OpenStudy (nubeer):

well i dont think its right thats why asking

OpenStudy (anonymous):

i only solved for s. thats what i thought you meant. i solved for s using this: t3= (2s-12) * 3

OpenStudy (nubeer):

its multiply with 3 not tthe power.

OpenStudy (anonymous):

show me your work, because you just confused me.

OpenStudy (nubeer):

multiply 3 with 2s then mltiply it with -12

OpenStudy (anonymous):

is it, t = 6s - 36?

OpenStudy (nubeer):

yes

OpenStudy (anonymous):

thats the answer?

OpenStudy (nubeer):

no .. it's t3.. u have to add t1,t2,t3

OpenStudy (anonymous):

so t3 = 6s - 36? and you have to add these three together?: 5s + 2s + 6s - 36 = 7s + 6s - 36 = 1s - 36

OpenStudy (anonymous):

i mean 14s - 36.

OpenStudy (nubeer):

yeah..

OpenStudy (anonymous):

so 14s - 36 is the answer? or is it, t = 14s - 36?

OpenStudy (nubeer):

its t=

OpenStudy (anonymous):

can you help me with just one more problem?

OpenStudy (nubeer):

depends on the question

OpenStudy (anonymous):

A restaurant has 45 tables. Each table seats 4 people. The manager has 4 employees to prepare the tables. The chart shows how fast each employee works. Use this information answer the following questions. A. In order to get the tables ready in the least amount of time, which two employees should the manager ask to fold napkins? Which two employees should the manager ask to set the tables? B. If everyone starts working at the same time, how quickly can the employees get the tables ready for dinner? The first pair of employees to finish their job should help the other pair to finish the tables. Explain the answer.

OpenStudy (nubeer):

lol way too long to solve.. sorry i will be going now..

OpenStudy (anonymous):

but its easier than the previous question, i just need a little help.

OpenStudy (nubeer):

well then what exactly u need little help with/

OpenStudy (anonymous):

Just A and B. Read it. It's easy for you.

OpenStudy (anonymous):

I just need help with Parts A and B only. Please?

OpenStudy (nubeer):

u need to take pic again because i see the picture inverted here

OpenStudy (nubeer):

stacie and nick should do the napkins and the other 2 should do the table.

OpenStudy (anonymous):

Okay, thanks. Can you help me with part B please? @nubeer